# Singular Bayesian Neural Networks

### v1

@@ E.10.3 Approach B: Modified Function Class with High-Probability Inclusion
    $\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r)$ by construction, where $C_{i}(\delta)=C_{i}^{A}(\delta)C_{i}^{B}(\delta)$ (all functions satisfy the spectral norm and rank constraints).
@@
    The BNN predictor restricted to $\mathcal{E}_{\text{good}}$ lies in $\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta))}$ by the same argument as Lemma E.11.
@@
By the same reasoning as Approach A:
$\mathbb{E}[\ell(f_{\mathrm{BNN}}(x),y))]$ $\leq\frac{1}{m}\sum_{i=1}^{m}\ell(f_{\mathrm{BNN}}(x_{i}),y_{i})$
@@
Both approaches yield the same final bound structure, with an additive $\delta$ term representing the high-probability relaxation.
  • Approach A is conceptually cleaner: it decomposes the predictor into good and bad events.
@@ E.11 Practical Enforcement Remarks
Alternatively, one may retain the high-probability statement and use Approaches A or B from Section E.10 to carry an additional failure probability $\delta_{\text{total}}$ through the final generalization bound. This is often more practical than truncation, as it requires no modification to the variational family.
@@
  • If using unbounded variational families without projection, apply the high-probability analysis from Section E.10 with appropriately chosen $\delta_{\text{total}}$.

### v2

@@ E.10.3 Approach B: Modified Function Class with High-Probability Inclusion
    $\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta)\subseteq\mathcal{F}_{D}^{\mathrm{Pinto}}(C(\delta),r)$ by construction, where $C_{i}(\delta)=C_{i}^{A}(\delta)C_{i}^{B}(\delta)$.
@@
    $f_{\mathrm{BNN}}^{\mathrm{good}}(\cdot;\theta)\in\overline{\mathrm{conv}(\mathcal{F}_{D}^{\mathrm{hp}}(\theta,\delta))}$ by the same argument as Lemma E.11, applied to the conditional posterior $q(\cdot\mid\mathcal{E}_{\text{good}})$.
@@
By the same reasoning as Approach A—decomposing $f_{\mathrm{BNN}}$ as in (22), applying Assumption E.27 to obtain (29), and bounding the population loss of $f_{\mathrm{BNN}}^{\mathrm{good}}$ via Theorem E.15—we obtain:
$\mathbb{E}_{(x,y)}[\ell(f_{\mathrm{BNN}}(x),y)]$ $\leq\frac{1}{m}\sum_{j=1}^{m}\ell(f_{\mathrm{BNN}}^{\mathrm{good}}(x_{j}),y_{j})$
@@
Both approaches yield the same final bound structure, with an additive $\delta$ term from the high-probability relaxation and a convexity requirement on the loss (Assumption E.27).
  • Approach A is conceptually cleaner: it decomposes the predictor into good and bad components at the function level.
@@ E.11 Practical Enforcement Remarks
Alternatively, one may retain the high-probability statement and use Approaches A or B from Section E.10 to carry an additional failure probability $\delta_{\text{total}}$ through the final generalization bound. This is often more practical than truncation, as it requires no modification to the variational family. This approach additionally requires loss convexity (Assumption E.27).
@@
  • If using unbounded variational families without projection, apply the high-probability analysis from Section E.10 with appropriately chosen $\delta_{\text{total}}$ and ensure the loss function satisfies Assumption E.27.